# Pentagon relation in quantum cluster scattering diagrams

### v4

@@ 3.2.1. Finite type
Here we consider the case where {,}$\{\cdot,\cdot\}\neq 0$. As in the classical case [Nak23, §III.1.5], without loosing generality, we may assume that $\{e_{2},e_{1}\}=1$ and $\delta_{2}\geq\delta_{1}$. Thus, the initial exchange matrix $B$ in (2.1) is given by
@@ 3.2.2. Affine type
(a). Type $A_{1}^{(1)}$. Let $(\delta_{1},\delta_{2})=(2,2)$. Note that, for any primitive $n\in N_{\mathrm{pr}}^{+}$, we have $\delta(n)=2$. The following description of the CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [GHKK18, Rei09, Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
@@
Indeed, one can prove the relation (3.75) completely based on the pentagon relation by modifying the proof of [Mat21] for the classical case. The detail will be found in Appendix A. This is the simplest example of consistency relations involving infinite product. The formula appeared in [DGS11, Eq. (1.3)] without proof in an alternative setting of quantum dilogarithms. See also an alternative derivation in view of the quantum affine algebra of type $A_{1}^{(1)}$ [Sug22, Theorem 5.1].
(b). Type $A_{2}^{(2)}$. Let $(\delta_{1},\delta_{2})=(1,4)$. The situation is parallel to $A_{1}^{(1)}$ though a little more complicated. The following description of the CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
@@ 4.1. Principal $x$-representation
(4.2) $\{\tilde{n},\tilde{n}^{\prime}\}$ :=$:=\{n,n^{\prime}\}+\langle n^{\prime},m\rangle-\langle n,m^{\prime}\rangle,$
(4.3) $\{\tilde{m},\tilde{m}^{\prime}\}$ :=$:=\{n,n^{\prime}\}+\langle n^{\prime},m\rangle-\langle n,m^{\prime}\rangle,$
@@
(4.7) $\begin{matrix}\tilde{p}^{*}\colon&N&\rightarrow&\tilde{M}^{\circ}\\ &n&\mapsto&\{\cdot,n\}\end{matrix},\quad\{(n^{\prime},m^{\prime}),n\}=\{n^{\prime},n\}+\langle n,m^{\prime}\rangle.$
@@
(4.9) $-\{\tilde{p}^{*}(n),\tilde{m}\}=\{\tilde{m},\tilde{p}^{*}(n)\}=\langle n,\tilde{m}\rangle.$
@@
(4.10) $\begin{split}\{\tilde{m},\tilde{p}^{*}(n)\}&=\{(m^{\prime},n^{\prime}),(p^{*}(n),n)\}\\ &=\{n^{\prime},n\}+\langle n,m^{\prime}\rangle-\langle n^{\prime},p^{*}(n)\rangle\\ &=\langle n,m^{\prime}\rangle=\langle n,\tilde{m}\rangle.\end{split}$
@@
Let $x$ be a formal symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[x]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $x^{\tilde{m}}$ ($\tilde{m}\in\tilde{P})$ and the relations
(4.11) $x^{\tilde{m}}x^{\tilde{m}^{\prime}}=q^{-\{\tilde{m},\tilde{m}^{\prime}\}}x^{\tilde{m}+\tilde{m}^{\prime}}.$
@@
Remark 4.1.
[…]
(4.18) $-\Lambda=\begin{pmatrix}O&D\\ -D&DB\end{pmatrix},\quad D=\mathrm{diag}(\delta_{1}^{-1},\cdots,\delta_{r}^{-1}).$
Let $\tilde{B}$ be the principal extension of the matrix $B$,
(4.21) $\tilde{B}=\begin{pmatrix}B\\ I\end{pmatrix}.$
They satisfy the relation
(4.24) $-\Lambda\tilde{B}=\begin{pmatrix}D\\ O\end{pmatrix},$
which is equivalent to the duality relation (4.9). Such a pair $(\Lambda,\tilde{B})$ is called a compatible pair in [BZ05]. The matrix $\Lambda$ was used to define the $q$-commutative relation for $x^{\tilde{m}}$ in [BZ05]. The minus sign for $-\Lambda$ in (4.18) agrees with the minus sign for $-\{\tilde{m},\tilde{m}^{\prime}\}$ in the relation (4.11).
@@ 4.2. Action of dilogarithm elements
Observe that this is indeed the automorphism part of the Fock-Goncharov decomposition of mutations of the quantum $x$-variables with principal coefficients in [BZ05]. See also [Man21, DM21, CFMM20] for an alternative approach, where the same representation is described through the adjoint action of the quantum dilogarithm on $\mathcal{R}_{q}(x)$ as in Remark 2.6.
@@ 4.3. Reduction of wall elements in a QCSD
Let $\mathcal{S}_{q}(x)$ be the subset of $\mathcal{R}_{q}(x)$ where the coefficients $c_{\tilde{m}}$ in (4.12) are restricted in $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$. Then, $\mathcal{S}_{q}(x)$ is an algebra over $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$ by the same relation (4.11). The following lemma is a quantum analogue of the one in [GHKK18, Appendix C.3 Step II], [Nak23, Lemma III.5.8].
@@
A $\mathbb{Q}(q^{1/\delta_{0}})$-power of a quantum dilogarithm element $\Psi_{a,b}[n]^{c}$ $(c\in\mathbb{Q}(q^{1/\delta_{0}}))$ acts on $\mathcal{S}_{q}(x)$ if and only if $a=1/s\delta(n)$ for a divisor $s$ of $\delta_{0}/\delta(n)$ and $c\in\mathbb{Z}$.
@@
The if-part is true by (4.27). Let us prove the only-if-part. Assume that $\Psi_{a,b}[n]^{c}$ acts on $\mathcal{S}_{q}(x)$. Then, in the second expression of the RHS of (4.25), the term
(4.28) $c\frac{q^{2j\langle n,\tilde{m}\rangle}-1}{q^{2ja}-1}\frac{(-1)^{j+1}}{j}q^{ja}q^{jb}$
should at least belong to $\mathbb{Q}[q^{\pm 1/\delta_{0}}]$ for any $j$. Let $a=1/s\delta(n)$, where $s$ is a positive rational number. Then, the condition implies $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{P}$. By the assumption for $\tilde{P}$, $\tilde{P}$ contains a basis of $\tilde{M}^{\circ}$. It follows that $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{M}^{\circ}$. Due to the definition of $\delta(n)$, there is some $\tilde{m}\in\tilde{M}^{\circ}$ such that $\langle\delta(n)n,\tilde{m}\rangle=1$. This implies $s\in\mathbb{Z}$. On the other hand, let $t=\delta_{0}/\delta(n)\in\mathbb{Z}_{>0}$. Then, $a=t/s\delta_{0}$. Since we require that $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, $s$ is a divisor of $t$. Finally, by (4.27) and the assumption, we have $c\in\mathbb{Z}$. ∎
We have the following important corollary of Lemma 4.2.
@@
(4.29) $\Psi_{1/\delta(n),b}[n]^{\pm c}\quad(n\in N^{+},\ b\in(1/\delta_{0})\mathbb{Z},\ c\in\mathbb{Z}).$
@@
This can be proved inductively on the degree of $n$ as in the classical case by considering the consistency relations around the perpendicular joints. See Step 2 in the proof of [GHKK18, Prop. C.13]. For the reader’s convenience, we give the key point of the inductive step. Suppose that a scattering diagram have the walls whose wall elements have the form (4.29) with $\deg n\leq\ell$ and that, for a small loop around a perpendicular joint $\gamma$, the consistency relation around $\gamma$ holds modulo $G^{>\ell}$. By the if-part of Lemma 4.2, all wall elements act on $\mathcal{S}_{q}(x)$. Then, we may add walls whose wall elements have the form (4.29) so that the consistency relation around $\gamma$ holds modulo $G^{>\ell+1}$. ∎
Example 4.4.
All examples of QCSDs in Section 3 have the wall elements of the form (4.29). Moreover, there are no negative powers $\Psi_{1/\delta(n),b}[n]^{-1}$.
@@ 5. Application to positivity and nonpositivity of QCSDs
In the rest of the paper we apply the pentagon relation, together with the combinatorics of quantum data, to study the positivity and nonpositivity property of QCSDs.
@@ 5.1. Positivity of QCSDs
Let us briefly summarize known results concerning the positivity of CSDs and QCSDs. Below, “skew-symmetric” means the initial exchange matrix in (2.1) is skew-symmetric.
  1. (1)
    Positivity of theta functions of CSD [GHKK18]. For any CSD (including nonskew-symmetric one), nonzero coefficients of any theta function are positive integers (positivity). In addition, nonzero multiplicities of the product of elements of theta basis are positive integers (strong positivity).
  2. (2)
    Positivity of theta functions of QCSD [DM21]. For any skew-symmetric QCSD, nonzero coefficients of any (quantum) theta function belong to $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$ (positivity). In addition, nonzero multiplicities of the product of elements of theta basis belong to $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$ (strong positivity).
  3. (3)
    Positivity and nonpositivity of quantum greedy basis [LLRZ14]. Consider a cluster algebra of rank 2 with the initial exchange matrix
    Then, the positivity of the quantum greedy elements fails for $(\delta_{1},\delta_{2})\allowbreak=(2,3)$, $(2,5)$, $(3,4)$, $(4,6)$, for example. Also, it was conjectured [LLRZ14, Conj. 14] that the strong positivity holds for the quantum greedy basis if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$. It is known that in the classical case the greedy basis and the theta basis coincide [${}^{+}$17]. Also, the above nonpositivity of the quantum greedy elements was explained in view of the nonpositivity of the wall functions in the corresponding QCSD [CFMM20] for the case $(\delta_{1},\delta_{2})=(2,3)$.
We see in the above that there is an intriguing discrepancy of the positivity property between CSDs and QCSDs especially in the nonskew-symmetric case. Then, we have the following natural questions.
  1. Q1:
    Exactly when and why does the positivity property of a nonskew-symmetric QCSD hold or fail?
  2. Q2:
    When the positivity property fails for a QCSD, how is the positivity property restored in the classical limit $q\rightarrow 1$?
Based on various results by [GHKK18, Nak23, Man21, DM21, LLRZ14, ${}^{+}$17, CFMM20], and also in view of Proposition 4.3, we formulate the positivity of QCSD as follows:
@@ 2. Q2:
Namely, the negative powers $\Psi_{1/\delta(n),b}[n]^{-c}$ ($c\in\mathbb{Z}_{>0}$) do not appear as wall elements. We say that a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive if it is equivalent to a consistent scattering diagram that is a positive realization.
@@
All examples of QCSDs in Section 3 are positive realizations, thus, positive as observed in Example 4.4.
For a QCSD, a theta function is defined by broken lines [DM21] as in the classical one [GHKK18].
@@
Every theta function of a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ has a formal power series of the form
@@
if $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive.
This was proved in [DM21, Theorem 3.9] by the same method for the classical case [GHKK18, Theorem 1.13, Remark 3.2]. Similarly, the positivity of a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ also implies the strong positivity of the theta basis [DM21, Theorem 3.10] as in the classical case [GHKK18].
We also expect that the converse of Proposition 5.3 holds. Let us concentrate on a particular situation, which is relevant to the forthcoming Theorem 5.6.
Example 5.4.
Consider the rank 2 case. Assume that there is some outgoing wall $\mathbf{w}$ in $\mathfrak{D}_{\mathfrak{s}}^{q}$ whose normal vector is $n=(1,1)$ and whose wall element has the form $\Psi_{1/\delta(n),b}[n]^{-c}$ ($c\in\mathbb{Z}_{>0}$), and it is not cancelled by the same type of walls with positive power. Take a point $Q$ in the same quadrant with $\mathbf{w}$. Consider a theta function $\theta_{Q,\tilde{m}_{0}}$ to which a broken line $\gamma$ bending only at $\mathbf{w}$ contributes. Then, $\gamma$ may carry a coefficient $c_{\gamma}$ not in $\mathbb{Z}_{\geq 0}[q^{\pm 1/\delta_{0}}]$. Also, the contribution of $\gamma$ to $\theta_{Q,\tilde{m}_{0}}$ is not cancelled from any other broken lines. Indeed, if there is such a broken line $\gamma^{\prime}$, it should bend on both of the two incoming walls and ends at $Q$. Thus, $\gamma^{\prime}$ should encircle the origin completely. But, this is impossible due to the definition of broken lines. Therefore, the converse of Proposition 5.3 holds in this case.
As already mentioned, the following result is known.
@@
Every skew-symmetric QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive.
The proof of [DM21] is based on the Donaldson-Thomas theory on quiver-representations, which is applicable only to the skew-symmetric case. Below we study the positivity of the nonskew-symmetric case by means of the pentagon relation.

### v7

@@ 3.2.1. Finite type
Here we consider the case where {,}$\{\cdot,\cdot\}\neq 0$. Without loss of generality, we may assume that $\{e_{2},e_{1}\}=1$ and $\delta_{2}\geq\delta_{1}$. Thus, the initial exchange matrix $B$ in (2.1) is given by
@@
(b) $B_{2}$$\begin{bmatrix}1\\ 0\end{bmatrix}_{1}$$\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{2}}$$\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{2}}$$\begin{bmatrix}1\\ 2\end{bmatrix}_{1}$
(c) $G_{2}$$\begin{bmatrix}1\\ 0\end{bmatrix}_{1}$$\begin{bmatrix}0\\ 1\end{bmatrix}_{\frac{1}{3}}$$\begin{bmatrix}1\\ 1\end{bmatrix}_{\frac{1}{3}}$$\begin{bmatrix}2\\ 3\end{bmatrix}_{1}$$\begin{bmatrix}1\\ 2\end{bmatrix}_{\frac{1}{3}}$$\begin{bmatrix}1\\ 3\end{bmatrix}_{1}$
@@ 3.2.2. Affine type
(a). Type $A_{1}^{(1)}$. Let $(\delta_{1},\delta_{2})=(2,2)$. Note that, for any primitive $n\in N_{\mathrm{pr}}^{+}$, we have $\delta(n)=2$. The following description of a CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [GHKK18, Rei09, Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
@@
Indeed, one can prove the relation (3.25) completely based on the pentagon relation by modifying the proof of [Mat21] for the classical case. The detail will be found in Appendix A. This is the simplest example of consistency relations involving infinite products. The formula appeared in [DGS11, Eq. (1.3)] without proof in an alternative setting of quantum dilogarithms. See also an alternative derivation in view of the quantum affine algebra of type $A_{1}^{(1)}$ [Sug23, Theorem 5.1].
(b). Type $A_{2}^{(2)}$. Let $(\delta_{1},\delta_{2})=(1,4)$. The situation is in parallel with $A_{1}^{(1)}$ though a little more complicated. The following description of the CSD $\mathfrak{D}_{\mathfrak{s}}$ is known [Rea20, Nak23, Mat21]: The walls of $\mathfrak{D}_{\mathfrak{s}}$ are given by
@@ 4.1. Principal $x$-representation
(4.2) $\{\tilde{n},\tilde{n}^{\prime}\}_{\tilde{N}}$ :=$:=\{n,n^{\prime}\}+\langle n^{\prime},m\rangle-\langle n,m^{\prime}\rangle,$
(4.3) $\{\tilde{m},\tilde{m}^{\prime}\}_{\tilde{M}^{\circ}}$ :=$:=-\{n,n^{\prime}\}-\langle n^{\prime},m\rangle+\langle n,m^{\prime}\rangle,$
@@
(4.5) $\begin{matrix}\tilde{p}^{*}\colon&N&\rightarrow&\tilde{M}^{\circ}\\ &n&\mapsto&\{\cdot,n\}_{\tilde{N}}\end{matrix},\quad\{(n^{\prime},m^{\prime}),n\}_{\tilde{N}}=\{n^{\prime},n\}+\langle n,m^{\prime}\rangle.$
@@
(4.7) $\{\tilde{p}^{*}(n),\tilde{m}\}_{\tilde{M}^{\circ}}=\langle n,\tilde{m}\rangle.$
@@
(4.8) $\begin{split}\{\tilde{p}^{*}(n),\tilde{m}\}_{\tilde{M}^{\circ}}&=\{(p^{*}(n),n),(m^{\prime},n^{\prime})\}_{\tilde{M}^{\circ}}\\ &=-\{n,n^{\prime}\}-\langle n^{\prime},p^{*}(n)\rangle+\langle n,m^{\prime}\rangle\\ &=\langle n,m^{\prime}\rangle=\langle n,\tilde{m}\rangle.\end{split}$
Remark 4.1.
The bilinear form {,}$\{\cdot,\cdot\}_{\tilde{M}^{\circ}}$ in (4.3) has the representation matrix with respect to the basis $f_{1}$, …, $f_{r}$, $e_{1}$, …, $e_{r}$ of $\tilde{M}^{\circ}$ as
(4.9) $\Lambda:=\begin{pmatrix}O&-D\\ D&-DB\end{pmatrix},\quad D=\mathrm{diag}(\delta_{1}^{-1},\cdots,\delta_{r}^{-1}).$
Let $\tilde{B}$ be the principal extension of the matrix $B$,
(4.10) $\tilde{B}=\begin{pmatrix}B\\ I\end{pmatrix}.$
They satisfy the relation
(4.11) $-\Lambda\tilde{B}=\begin{pmatrix}D\\ O\end{pmatrix},$
which is equivalent to the duality relation (4.7). Such a pair $(\Lambda,\tilde{B})$ is called a compatible pair in [BZ05].
@@
  • (iii).
    $f_{1}$, …, $f_{r}\in\tilde{P}$.
For example, we take $\sigma$ as the cone generated by $f_{1}$, …, $f_{r}$, $\tilde{p}^{*}(e_{1})$, …, $\tilde{p}^{*}(e_{r})$, which are a basis of $\tilde{M}^{\circ}$. (In [GHKK18], the condition (iii) was not assumed.)
@@
Let $x$ be a formal symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[x]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $x^{\tilde{m}}$ ($\tilde{m}\in\tilde{P})$ and the $q$-commutative relations
(4.12) $x^{\tilde{m}}x^{\tilde{m}^{\prime}}=q^{\{\tilde{m},\tilde{m}^{\prime}\}_{\tilde{M}^{\circ}}}x^{\tilde{m}+\tilde{m}^{\prime}}.$
@@ 4.2. Action of dilogarithm elements
Observe that this is indeed the automorphism part of the Fock-Goncharov decomposition of mutations of the quantum $x$-variables with principal coefficients in [BZ05]. See also [Man21, DM21, CFMM24] for an alternative approach, where the same representation is described through the adjoint action of the quantum dilogarithm on $\mathcal{R}_{q}(x)$ as in Remark 2.6.
@@ 4.3. Reduction of wall elements in a QCSD
Let $\mathcal{S}_{q}(x)$ be the subset of $\mathcal{R}_{q}(x)$ consisting of the elements such that the coefficients $c_{\tilde{m}}$ in (4.13) are in $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$. Then, $\mathcal{S}_{q}(x)$ is an algebra over $\mathbb{Z}[q^{\pm 1/\delta_{0}}]$ by the same relation (4.12). The following lemma is a quantum analog of the one in [GHKK18, Appendix C.3 Step II], [Nak23, Lemma III.5.8].
@@
A $\mathbb{Q}(q^{1/\delta_{0}})$-power of a quantum dilogarithm element $\Psi_{a,b}[n]^{c}$ $(c\in\mathbb{Q}(q^{1/\delta_{0}}))$ acts on $\mathcal{S}_{q}(x)$ if and only if $a=1/s\delta(n)$ for a divisor $s$ of $\delta_{0}/\delta(n)$ and $c\in\mathbb{Z}[q^{\pm 1/\delta_{0}}]$.
@@
First, we prove the if-part. By the assumption $c\in\mathbb{Z}[q^{\pm 1/\delta_{0}}]$, $\Psi_{1/s\delta(n),b}[n]^{c}$ is decomposed into a product of $\Psi_{1/s\delta(n),b_{i}}[n]^{c_{i}}$ ($b_{i}\in(1/\delta_{0})\mathbb{Z}$, $c_{i}\in\mathbb{Z}$). By (4.19), each factor acts on $\mathcal{S}_{q}(x)$. Thus, $\Psi_{a,b}[n]^{c}$ acts on $\mathcal{S}_{q}(x)$. Next, we prove the only-if-part. Assume that $\Psi_{a,b}[n]^{c}$ acts on $\mathcal{S}_{q}(x)$. Then, by (4.17), we have a necessary condition
(4.20) $c\frac{q^{2j\langle n,\tilde{m}\rangle}-1}{q^{2ja}-1}\frac{(-1)^{j+1}}{j}q^{ja}q^{jb}\in\mathbb{Q}[q^{\pm 1/\delta_{0}}]\quad(j\in\mathbb{Z}_{>0}).$
Let $a=1/s\delta(n)$, where $s$ is a positive rational number. Then, (4.20) implies the condition $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{P}$. By the assumption for $\tilde{P}$, $\tilde{P}$ contains a basis of $\tilde{M}^{\circ}$. It follows that $s\langle\delta(n)n,\tilde{m}\rangle\in\mathbb{Z}$ for any $\tilde{m}\in\tilde{M}^{\circ}$. Due to the definition of $\delta(n)$, there is some $\tilde{m}\in\tilde{M}^{\circ}$ such that $\langle\delta(n)n,\tilde{m}\rangle=1$. Thus, we have $s\in\mathbb{Z}$. On the other hand, let $t=\delta_{0}/\delta(n)\in\mathbb{Z}_{>0}$. Then, $a=t/s\delta_{0}$. Since we require that $a\in(1/\delta_{0})\mathbb{Z}_{>0}$, $s$ is a divisor of $t$. This finishes the condition for $a$. Meanwhile, the condition (4.20) also implies $c\in\mathbb{Q}[q^{\pm 1/\delta_{0}}]$. Then, $\Psi_{a,b}[n]^{c}$ is decomposed into a product of $\Psi_{1/s\delta(n),b_{i}}[n]^{c_{i}}$ ($b_{i}\in(1/\delta_{0})\mathbb{Z}$, $c_{i}\in\mathbb{Q}$). By (4.19), we have $c_{i}\in\mathbb{Z}$. Thus, we have $c\in\mathbb{Z}[q^{\pm 1/\delta_{0}}]$. ∎
Below, we only use the if-part of Lemma 4.2.
@@
(4.21) $\Psi_{1/\delta(n),b}[n]^{c}\quad(n\in N^{+},\ b\in(1/\delta_{0})\mathbb{Z},\ c\in\mathbb{Z}).$
@@
We follow Step 2 in the proof of [GHKK18, Prop. C.13]. This is proved inductively on the degree of $n$, by considering the consistency relations around the perpendicular joints. Here, we only describe the key point of the inductive step. Suppose that a scattering diagram $\mathfrak{D}_{\ell}$ has the walls with wall elements of the form (4.21) with $\deg n\leq\ell$ and that, for a small loop $\gamma$ around a perpendicular joint $\mathfrak{j}$, the consistency relation $\mathfrak{p}_{\gamma,\mathfrak{D}_{\ell}}=\mathrm{id}$ holds modulo $G^{>\ell}$. By Lemma 4.2, $\mathfrak{p}_{\gamma,\mathfrak{D}_{\ell}}$ acts on $\mathcal{S}_{q}(x)$. By adding walls around $\mathfrak{j}$ with wall elements $\Psi_{a,b}[n]^{c}$ ($\deg n=\ell+1$) to $\mathfrak{D}_{\ell}$, we have a scattering diagram $\mathfrak{D}_{\ell+1}$ such that $\mathfrak{p}_{\gamma,\mathfrak{D}_{\ell+1}}=\mathrm{id}$ modulo $G^{>\ell+1}$. Since we consider the action of $\Psi_{a,b}[n]^{c}$ modulo $G^{>\ell+1}$, the second expression of (4.17) is simplified as
(4.22) $x^{\tilde{m}}\biggl{(}1+c\frac{q^{2\langle n,\tilde{m}\rangle}-1}{q^{2a}-1}q^{a}q^{b}x^{\tilde{p}^{*}(n)}\biggr{)}.$
We may further concentrate on $\tilde{m}$ with $\langle\delta(n)n,\tilde{m}\rangle=1$, because other cases are propotional to it with factors that are independent of $\Psi_{a,b}[n]^{c}$. In particular, the action of $\Psi_{1/\delta(n),0}[n]^{c}$ on such $\tilde{m}$ is given by
(4.23) $x^{\tilde{m}}(1+cq^{a}x^{\tilde{p}^{*}(n)}).$
This means that, for each $n\in N^{+}$ ($\deg n=\ell+1$), the (possibly multiple) added wall elements $\Psi_{a,b_{i}}[n]^{c_{i}}$ ($i=1$, 2, …) in $\mathfrak{D}_{\ell+1}$ is replaced with a single wall element $\Psi_{1/\delta(n),0}[n]^{c}$; moreover, we have $c\in\mathbb{Z}[q^{\pm 1/\delta_{0}}]$ by the induction assumption. Then, it is further decomposed into a product of $\Psi_{1/\delta(n),b}[n]^{c}$’s ($c\in\mathbb{Z}$). ∎
@@ 5. Application to positivity and nonpositivity of QCSDs
In the rest of the paper we apply the pentagon relation, together with the combinatorics of quantum data, to study the positivity and nonpositivity of QCSDs.
@@ 5.1. Positivity of QCSDs
Let us briefly summarize known results concerning the positivity and nonpositivity of CSDs and QCSDs.
Let us introduce the following notion.
Definition 5.1.
A CSD $\mathfrak{D}_{\mathfrak{s}}$ is a positive realization if the wall element of any wall has the form
(5.1) $\Psi[n]^{c\delta(n)}\quad(n\in N^{+},\ b\in(1/\delta_{0})\mathbb{Z},\ c\in\mathbb{Z}_{>0}).$
We say that a CSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive if it is equivalent to a CSD that is a positive realization.
The following fact is one of the main results of [GHKK18].
Theorem 5.2 ([GHKK18, Theorem 1.13]).
Every CSD $\mathfrak{D}_{\mathfrak{s}}$ is positive.
For each CSD, a theta function is defined by broken lines [GHKK18]. The definition of a broken line and Theorem 5.2 immediately imply the following positivity result.
Corollary 5.3 (Positivity of theta functions [GHKK18]).
For any CSD $\mathfrak{D}_{\mathfrak{s}}$, every nonzero coefficients of any theta function is a positive integer.
Based on Proposition 4.3, we introduced a parallel notion for QCSDs.
@@
We say that a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive if it is equivalent to a QCSD that is a positive realization.
@@
All examples of QCSDs in Section 3 are positive realizations.
A quantum theta function for a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is defined by broken lines [DM21] in the same way as in the classical one [GHKK18]. The following fact follows from the definition of a broken line.
@@
Suppose that $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive. Then, every quantum theta function of a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ has a formal power series of the form
@@
Observing the parallelism between the CSDs and QCSDs so far, it is natural to expect that Theorem 5.2 also holds for QCSDs. However, this is not true in general. We say that a QCSD is skew-symmetric if its initial exchange matrix $B$ is skew-symmetric. The following positivity result is known.
@@
Any skew-symmetric QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is positive.
The proof of [DM21] is based on the Donaldson-Thomas theory on quiver-representations, which applies exclusively to the skew-symmetric case.
On the other hand, the following example suggests the abundance of the nonpositive QCSDs.
Example 5.8 ([LLRZ14, CFMM24]).
Consider a cluster algebra of rank 2 with the initial exchange matrix
The (classical) greedy elements were introduced in [LLZ12]. Later they were identified with the theta functions [CGM+17]. Meanwhile, the quantum greedy elements were studied in [LLRZ14]. It turned out that the positivity of the coefficients of the quantum greedy elements fails for $(\delta_{1},\delta_{2})\allowbreak=(2,3)$, $(2,5)$, $(3,4)$, $(4,6)$, for example. Motivated by this result, the nonpositivity of quantum theta functions for $(\delta_{1},\delta_{2})=(2,3)$ was shown by [CFMM24]. Thus, the corresponding QCSD is nonpositive due to Proposition 5.6. Also, it was conjectured [LLRZ14, Conj. 14] that the positivity holds for the quantum greedy basis if $\delta_{1}|\delta_{2}$ or $\delta_{2}|\delta_{1}$.
Thus, there is an intriguing discrepancy of the positivity between CSDs and QCSDs especially in the nonskew-symmetric case. Then, we have the following natural questions.
  1. Q1:
    Exactly when and why does the positivity of a nonskew-symmetric QCSD hold or fail?
  2. Q2:
    When the positivity fails for a QCSD, how is the positivity restored in the classical limit $q\rightarrow 1$?
Below we study the above problems for the nonskew-symmetric QCSDs of rank 2 using the pentagon relation. Before that, let us present one immediate and important consequence of the positivity of QCSDs. We say that a CSD $\mathfrak{D}_{\mathfrak{s}}$ is with minimal support if $\mathrm{Supp}(\mathfrak{D}_{\mathfrak{s}})$ is minimal among all CSDs that are equivalent to $\mathfrak{D}_{\mathfrak{s}}$. If $\mathfrak{D}_{\mathfrak{s}}$ is a positive realization, then $\mathfrak{D}_{\mathfrak{s}}$ is with minimal support because it is impossible to cancel the wall elements of any walls due to the positivity.
Proposition 5.9.
Suppose that a QCSD $\mathfrak{D}_{\mathfrak{s}}^{q}$ is a positive realization. Then, $\mathrm{Supp}(\mathfrak{D}_{\mathfrak{s}}^{q})$ coincides with $\mathrm{Supp}(\mathfrak{D}_{\mathfrak{s}})$, where $\mathfrak{D}_{\mathfrak{s}}$ is a CSD with minimal support.
Proof.
By replacing each wall element $\Psi_{1/\delta(n),b}[n]^{c}$ in $\mathfrak{D}_{\mathfrak{s}}^{q}$ with $\Psi[n]^{c\delta(n)}$, we obtain a CSD $\mathfrak{D}_{\mathfrak{s}}$, which is a positive realization. ∎